# On the relationship between the Wasserstein distance and differences in life expectancy at birth

### v1

@@ 1 Introduction
where $\mathcal{J}(P,Q)$ denotes a set of all couplings (also called transport plans) between $P$ and $Q$. The $\inf_{J\in\mathcal{J}(P,Q)}$ denotes the infimum, i.e., among all transport plans, we are looking for the one with the smallest cost. For each pair, $(x,y)$, the cost of moving mass from $x$ to $y$ is given by $\|x-y\|^{p}$. Accordingly, $\|x-y\|^{p}$ defines how the distance between $x$ and $y$ is measured. If we set $p=1$, we are considering the average absolute distance moved. This is usually called the Earth Mover’s distance (Wasserman 2019).
Further, it has been shown that in a special case - when considering distributions from a one-dimensional space and using the absolute difference, $p=1$, as the cost function - the Wasserstein distance simplifies to,
$W_{1}(P,Q)=\int_{-\infty}^{+\infty}|F-G|,$ (2)
where $F$ and $G$ are the respective cumulative distribution functions (CDFs) of $P$ and $Q$. Please refer to Santambrogio (2015), chapter 2 for more details on the equivalence of both equations.
@@ 3.1 Sampling from the Human Mortality Database
The database includes period life tables for several countries in various years (sometimes going back to the 1800s). Each period life table is available for women, men, and for both sexes combined. It is computationally demanding to compare differences in $e_{0}$ and $W_{1}$ Wasserstein distance for all possible pairs in the dataset. For this reason, we focus on life tables for both sexes combined and sample 5 000 randomly selected country pairs from the subset. The cross-country comparison refers to one randomly select year, i.e., we do not compare period life tables over time. In addition to $W_{1}$ and differences in $e_{0}$, we calculate the non-overlap index - also called Jaccard index - (Shi et al. 2022) and the Kullback-Leibler divergence for each pair. This allows comparing the $W_{1}$ Wasserstein result to two alternatives measures for distributional similarity. One of the 5 000 sampled pairs is Denmark and Belgium in 1887. For this pair, the difference in $e_{0}$ equals the $W_{1}$ Wasserstein distance ($3.33$). The non-overlap index is $0.14$, while the Kullback-Leibler divergence is $0.02$. The results for all 5 000 pairs is depicted in Figure 1(b) and 2.
@@
In most cases, a large $e_{0}$ difference coincides with a large $W_{1}$ Wasserstein distances. However, we do find cases where the gap in $e_{0}$ is very small but $W_{1}$ Wasserstein suggests large distributional differences. This is observed, for example, when comparing England & Wales with Iceland in 1849. Both countries show a $e_{0}$ value of about 37.3. Yet, the age-at-death distributions differ substantially (see Figure 4(b)). While Iceland shows higher infant mortality as compared to England & Wales, mortality is lower for Iceland at older ages. Accordingly, the net difference in survivorship over age is small but there are still large absolute differences in $l(x)$ (see 4(a)).

### v2

@@ 1 Introduction
where $\mathcal{J}(P,Q)$ denotes a set of all couplings (also called transport plans) between $P$ and $Q$. The $\inf_{J\in\mathcal{J}(P,Q)}$ denotes the infimum, i.e., among all transport plans, we are looking for the one with the smallest cost. For each pair, $(x,y)$, the cost of moving mass from $x$ to $y$ is given by $\|x-y\|^{p}$. Accordingly, $\|x-y\|^{p}$ defines how the distance between $x$ and $y$ is measured. The outer exponent $1/p$ ensures that the resulting distance has the same units as the underlying metric.
When $p=1$ the measure is called the Earth Mover’s Distance (EMD) and represents the minimum average absolute distance you must move mass to transform $P$ into $Q$ (Wasserman 2019).
The equation 1 gives a general formulation of the Wasserstein distance. In the one-dimensional case, the optimal transport problem is much simpler. Let $F_{P}$ and $F_{Q}$ be the cumulative distribution functions (CDF) of $P$ and $Q$, respectively. Then, the inverse of $F_{P}$ and $F_{Q}$ give the quantile functions $F_{P}^{-1}$ and $F_{Q}^{-1}$. The Wasserstein distance in the one-dimensional case, can be expressed as,
$W_{p}(P,Q)=\left(\int_{0}^{1}\left|F_{P}^{-1}(x)-F_{Q}^{-1}(x)\right|^{p}\,dx\right)^{1/p}.$ (2)
Further, it has been shown that in a special case - when considering distributions from a one-dimensional space and using the absolute difference, $p=1$, as the cost function, the Wasserstein distance simplifies to,
$W_{1}(P,Q)=\int_{-\infty}^{+\infty}|F_{P}(x)-F_{Q}(x)|dx.$ (3)
Please refer to Santambrogio (2015), chapter 2 for more details on the equivalence of both equations.
The equation 3 is key for assessing distributional differences between two life table age-at-death distributions. Using $p=1$ is appropriate when comparing age-at-death distributions because it treats transport costs linearly, i.e., moving probability mass from age ten to age zero indicates a cost equal to the distance between the two ages (|xy|=|100|=10)(|x-y|=|10-0|=10). In other words, with $p=1$, the distance is measured simply as the absolute difference in ages. If we were to use $p=2$ instead, the cost of transporting mass grows quadratically since the distance term becomes $(|x-y|^{2}=(|10-0|)^{2}=100)$. This quadratic scaling makes transports over long distances disproportionately expensive, e.g., between very old ages and very young ages. In contrast, the linear $p=1$ formulation preserves a direct interpretation of transport distance as the absolute age difference, which is often more meaningful for comparing age-at-death distributions.
@@ 2.3 Wasserstein distance and the difference in life expectancy at birth
Moreover, equation 10 indicates clearly the range of possible Wasserstein distance values. In case of two identical $l(x)$ functions, Wasserstein distance equals zero. If there is no overlap between the two $l(x)$ functions, Wasserstein distance reaches its maximum value. In a life table with single ages from age zero to age 100, the maximum Wasserstein distance value is therefore 100. The one-dimensional Wasserstein distance value can be interpreted in terms of the average age shift that is required to transform one distribution into another. If the distributions differ mainly because one population dies younger, Wasserstein distance reflects the magnitude of the age shift. Further, if the distributions differ only in shape (one more concentrated vs. one more spread out), Wasserstein distance captures the minimum mass movements required to reconcile these shapes.
Figure 1: Toy example for demonstrating the idea behind the Wasserstein distance using all-cause specific mortality
To make both distributions look the same, we need to move $0.6$ mass from age $70$ to age $10$. Hence, the Wasserstein distance is given by $0.6\cdot(|70-10|)=10$.
2.4 Using the Wasserstein distance for cause-specific mortality comparisons
Multi-decrement life tables allow the derivation of age-at-death distributions by specific causes of death (Preston et al. 2001). As before, we set the life table radix $l_{0}=1$. Summing $d(x,c)$ over all ages $x$ and causes $c$ therefore equals one, ensuring that the resulting cause-specific age-at-death distribution is a proper probability distribution. The Wasserstein distance for the one-dimensional all-cause distribution can be computed analytically using equation 10. In contrast, the two-dimensional Wasserstein distance does not have a closed-form solution. To obtain the optimal transport plan, we can use the Python Optimal Transport (POT) package, which implements the network simplex algorithm to solve the transport problem.
The two-dimensional Wasserstein distance requires specifying a cost matrix that determines how costly it is to move probability mass between age-cause combinations. As described above, we set the cost of moving mass across ages as the absolute age difference. Choosing the cost of movement between causes of death, however, is less straightforward. Setting the cross-cause cost to zero makes movements between causes free, i.e., the resulting Wasserstein distance would then ignore cause-specific differences and measure only the mismatch in age patterns between two distributions. At the opposite extreme, setting the cost to infinity makes movements between causes impossible. In this case, the algorithm solves the optimal transport problem for each cause separately. Intermediate choices are also possible. For example, setting the cause-movement cost equal to one implies that shifting mass by one year of age has the same cost as reallocating mass across causes. The appropriate scale therefore depends on the substantive interpretation of cause-of-death differences. The choice of the cost for moving mass determines how large Wasserstein distance can be. Mathematically, we can define the cost matrix for the two-dimensional Wasserstein distance as,
$M_{ij}=|x_{i}-x_{j}\ |+\lambda\mathbf{1}\{c_{i}\neq c_{j}\ \},$ (12)
where $|x_{i}-x_{j}|$ is the transport cost for moving mass between ages (measured in years), while $\lambda\mathbf{1}\{c_{i}\neq c_{j}\}$ is an indicator equal to one if the causes differ. The two-dimensional Wasserstein distance does not have a fixed maximum value because it depends on the size of $\lambda$. A pragmatic strategy is to anchor the cost of moving mass between causes to the all-cause (one-dimensional) Wasserstein distance. This can be done dynamically, i.e., assigning each pair of distributions its own cause-movement cost based on their one-dimensional Wasserstein distance, or statically, using a fixed all-cause Wasserstein distance value for all two-dimensional calculations. If we do not set the cost for moving mass between causes of death to infinity, the two-dimensional Wasserstein distance quantifies how much work is needed to align both age patterns and cause-of-death patterns simultaneously, where work is defined by the amount of death probability mass that must be shifted multiplied by the distance (or cost) over which it is moved.
Figure 2: Toy example for demonstrating the idea behind theWasserstein distance using cause-specific mortality
To make both distributions look the same, we first need to move $0.8$ mass from age $85$ to $80$ $(|80-85|\cdot 0.8=4)$. Then, we need to move $0.2$ mass between causes two and three (the age dimension is already equal for both distributions). The key question is how much does moving mass between two causes cost? If we use a $\lambda$ of $10$, the transport is quantified by $|60-60|+10\cdot 0.2=2$. Adding both transports together gives a two-dimensional Wasserstein distance value of six $(4+2=6)$.
@@ 3.1 Sampling from the Human Mortality Database
The database includes period life tables for several countries in various years (sometimes going back to the 1800s). Each period life table is available for women, men, and for both sexes combined. It is computationally demanding to compare differences in $e_{0}$ and $W_{1}$ Wasserstein distance for all possible pairs in the dataset. For this reason, we focus on life tables for both sexes combined and sample 5 000 randomly selected country pairs from the subset. The cross-country comparison refers to one randomly select year, i.e., we do not compare period life tables over time. In addition to $W_{1}$ and differences in $e_{0}$, we calculate the non-overlap index - also called Jaccard index - (Shi et al. 2022) and the Kullback-Leibler divergence for each pair. This allows comparing the $W_{1}$ Wasserstein result to two alternatives measures for distributional similarity. One of the 5 000 sampled pairs is Denmark and Belgium in 1887. For this pair, the difference in $e_{0}$ equals the $W_{1}$ Wasserstein distance ($3.33$). The non-overlap index is $0.14$, while the Kullback-Leibler divergence is $0.02$. The results for all 5 000 pairs is depicted in Figure 3(b) and 4.
@@
In most cases, a large $e_{0}$ difference coincides with a large $W_{1}$ Wasserstein distances. However, we do find cases where the gap in $e_{0}$ is very small but $W_{1}$ Wasserstein suggests large distributional differences. This is observed, for example, when comparing England & Wales with Iceland in 1849. Both countries show a $e_{0}$ value of about 37.3. Yet, the age-at-death distributions differ substantially (see Figure 6(b)). While Iceland shows higher infant mortality as compared to England & Wales, mortality is lower for Iceland at older ages. Accordingly, the net difference in survivorship over age is small but there are still large absolute differences in $l(x)$ (see 6(a)).
@@ 4 Conclusion
It also possible to use the Wasserstein distance in a two-dimensional setting. This requires the use of an algorithm to solve the optimal transport problem as there is no closed-form solution. Also, it is difficult to find a meaningful $\lambda$ parameter, i.e., the parameter defining the cost for moving mass between causes of death.